# The Minimal Attached Eddy in Wall Turbulence: Statistical Foundations, Inverse Identification and Influence Kernels

### v4

@@ 2.2  A simplified influence-function model and the log law
$U(y)$ $=\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\int_{h_{min}}^{h_{max}}\frac{I_{1}\left({y}/{h}\right)}{h}dh+U_{ref}$
$=\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\int_{y/h_{max}}^{y/h_{min}}\frac{I_{1}\left({y}/{h}\right)}{y/h}d(y/h)+U_{ref}$
$\approx\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\left[\int_{y/h_{max}}^{1}\frac{-a_{0}}{y/h}d(y/h)+\int_{1}^{a_{1}/a_{2}}\frac{a_{1}-a_{2}y/h}{y/h}d(y/h)\right]+U_{ref}$
@@
3  Predictions
@@
It is remarked that this eddy corresponds to a closed loop in the sense that the leg of the eddy aligned with the wall is canceled by an equal and opposite image vortex pair, and is thus consistent with the vorticity kinematics. Using the Biot–Savart law to compute the induced velocity $u_{1}$ in Eqn. 4, Figure 6 shows that such a simple attached eddy can reproduce the mean streamwise velocity accurately. Further, the second moments are also seen to be well-predicted as shown in Figure 6.
@@
Figure 6: Velocity moments and eddy influence function. Lines: Optimal eddy influence function. Symbols: Square hairpin at $\theta=60^{\circ}$. For the right figures, colors follow Figure 2.
@@
It is noted that an extensive optimization was performed on different eddy shapes (e.g. triangular hairpin) and even packet configurations. Yet, these did not yield improved predictions. The hairpin at an angle of $\theta=68^{\circ}$ appears to be most optimal. We now further evaluate the predictive capabilities of this model by using the $\theta=60^{\circ}$ configuration and comparing it to the experimental measurements from samie2018fully. As shown in Figures 7,8,9, the model remains remarkably accurate in predicting mean velocity, streamwise Reynolds stress and spectra over a range of Reynolds numbers.
@@
Figure 7: Mean velocity prediction using square hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols :  samie2018fully
@@
Figure 8: Streamwise velocity fluctuation predictions using square hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols :  samie2018fully
@@
Figure 9: Pre-multiplied streamwise energy spectra predictions at $y^{+}=1184$ using square hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols :  samie2018fully
4  Physical and Statistical Insights
It is useful to separate (i) the population law $(\beta,p(h),h_{\min},h_{\max})$ from (ii) the single-eddy template (geometry, inclination angle $\theta$, circulation scaling, and vortex-core regularization). In this study we adopt Townsend/Perry scaling $p(h)\propto h^{-3}$ and determine $\beta$ from the log-law relation (Eq. 21) using $\kappa=0.384$, after which $\beta$ is held fixed across the Reynolds-number comparisons in Section 3. The template is likewise fixed once (square hairpin with $\theta=60^{\circ}$ and a small Rankine core), so the subsequent moment and spectral comparisons should be interpreted as out-of-sample tests of that single-eddy choice rather than repeated parameter fitting. Under these conventions, the only continuously adjusted parameter used to anchor the data is the population amplitude $\beta$.
@@ 4  Physical and Statistical Insights
The inverse problem strongly constrains what any viable attached eddy must do in a wall-parallel plane. In particular, the inferred mean influence function $I_{1}(y/h)$ exhibits (a) a near-constant region for $y/h\lesssim 1$ and (b) rapid decay for $y/h\gtrsim 1$, implying that eddies larger than the observation height dominate the log-region population integral. The square hairpin + image realizes these features because the two legs and head generate partially canceling far-field contributions, while the inviscid image enforces wall compatibility and suppresses the leading-order $1/\rho^{2}$ tail that would otherwise make the planar influence integrals marginal (Appendix A). By contrast, a triangular hairpin produces a more monotone “ramp” in $I_{1}$ and does not reproduce the inferred plateau as cleanly (Figure 6). Experimentation with a pack of triangular hairpins with variable circulation strengths did not improve results beyond the square hairpin which appears optimal.
@@
4.1  Similarity structure and the Influence kernel
@@
$u_{h}(x,z;y)=\mathcal{U}\!\left(\frac{x}{h},\frac{z}{h},\frac{y}{h}\right)=\mathcal{U}(\xi,\zeta,\eta),$
where $\mathcal{U}$ is the unit-eddy footprint expressed in dimensionless coordinates. Taking the Fourier transform and changing variables yields the exact scaling $\widehat{u}_{h}(k_{x},k_{z};y)=\,h^{2}\,\widehat{\mathcal{U}}(\kappa_{x},\kappa_{z};\eta),$ where $\widehat{\mathcal{U}}(\kappa_{x},\kappa_{z};\eta)=\int_{\mathbb{R}^{2}}\mathcal{U}(\xi,\zeta,\eta)\,e^{-i(\kappa_{x}\xi+\kappa_{z}\zeta)}\,d\xi\,d\zeta.$ Therefore
@@
where the Influence kernel is defined by
@@
This identity is a concrete mathematical object that will help decode the energy spectrum in greater detail.
@@
Figure 10: Streamwise energy Influence kernel $I_{\phi}(\kappa_{x},\eta)$ for square hairpin at $\theta=60^{\circ}$. This figure highlights that the kernel has a broad, elevated region at small $\eta$ and small-to-moderate $\kappa$, and a more localized hump as $\eta$ increases. Also shown are contributions to $\phi_{xx}^{+}(k_{x}^{+})$ for $k_{x}^{+}=1\times 10^{-4}$ (magenta) and $k_{x}^{+}=1\times 10^{-3}$ (red) at $y^{+}=1184$ for $Re_{\tau}=20000$.
@@
The kernel $I_{\phi}(\kappa_{x},\eta)$ is a function of the dimensionless wavenumber $\kappa_{x}=k_{x}h$ at fixed relative height $\eta=y/h$ and is shown in Figure 10. It is immediately clear that eddies of scale $h<y^{+}$ have negligible influence on the statistics.
@@
For any fixed $\eta$, typical scaling arguments yield a concentration the influence $S(\kappa_{x},\eta)$ around $\kappa_{x}=O(1)$ because $\kappa_{x}$ is the only dimensionless streamwise wavenumber available to the unit eddy. Therefore, for a given size $h$ and fixed $y$, the dominant contribution of that eddy appears around
@@
equivalently $\lambda_{x}=2\pi/k_{x}=O(h)$. This explains the peak of the premultiplied spectra in Fig. 9. Additionally, the square hairpin Attached Eddy model is able to recover a near-perfect linear behavior of the streamwise energy spectrum in the intermediate scale range as shown in Fig. 11. Below, we will give more concrete and quantitative explanations.
@@ 4.2  From single-eddy spectra to the attached-eddy spectrum
The $1/k_{x}$ scaling in the logarithmic region arises when $F(k_{x};y)$ varies slowly with $k_{x}$ over an intermediate band. This occurs when the $\kappa$-window $k_{x}h_{\min},k_{x}h_{\max}]$ covers the portion of the kernel $S(\kappa,\eta)$ that carries most of the contribution, and when the effective $\eta$ values encountered (through $\eta=y/h=yk_{x}/\kappa$) lie in the log-layer-relevant range. In that band,
@@
The above $1/k_{x}$” argument corresponds to the case that $F(k_{x};y)$ is approximately constant over an intermediate $k_{x}$ range. However the low-$k_{x}$ behavior is entirely controlled by how $F(k_{x};y)$ varies as $k_{x}\to 0$. There is no universality guarantee there: when $k_{x}$ is small, the $\kappa$-window and the $\eta$-trajectory can force the integral to sample the least robust part of the kernel. In other words, for small $k_{x}$, this trajectory spends substantial “arc length” in the region of small $\kappa$ and small $\eta$.
@@
Physically, low $k_{x}$ corresponds to very long streamwise wavelengths (VLSM-like content). In an attached-eddy framework with independent marks (Poisson placement in the wall-parallel plane), the low-$k_{x}$ variance is built by incoherently summing contributions of the largest eddies. In a Poisson superposition, long-wavelength modes are fed by many large eddies with essentially random phases, so variances add without cancellation. Real wall turbulence exhibits organization, intermittency, and correlations at the largest scales; those effects can reduce energy in the very lowest modes relative to an independent-eddy superposition.  baars2020data1; baars2020data2 offer more quantitative insights in the form of 3 different eddy types, and that attached eddies represent one of these types and contribute a large fraction of the energy only at a intermediate range of scales in the log layer. The latter point has been well argued by  perry1995wall. The independent-mark hypothesis therefore tends to overpopulate the lowest wavenumbers unless correlations are explicitly modeled or packaged into larger marks. It is possible that the influence kernel $I_{\phi}$ can be used in an inverse problem setting to ‘design’ compositions of prototypical eddies to model desired spectral properties.

### v5

@@ 2.2  A simplified influence-function model and the log law
$U(y)$ $=\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\int_{h_{\min}}^{h_{\max}}\frac{I_{1}\left({y}/{h}\right)}{h}dh+U_{ref}$
$\approx\frac{2\beta}{1/h_{\min}^{2}-1/h_{\max}^{2}}\left[\int_{y/h_{\max}}^{1}\frac{-a_{0}}{y/h}d(y/h)+\int_{1}^{a_{1}/a_{2}}\frac{a_{1}-a_{2}y/h}{y/h}d(y/h)\right]+U_{ref}$
@@
It is worth noting that Eq. (18) provides a direct route from measurable quantities ($\kappa$, $Re_{\tau}$) and the eddy population density $\beta$ to the plateau value $a_{0}$ of the influence function, or conversely, from the inferred $a_{0}$ and a prescribed $\beta$ to a prediction of $\kappa$. This connection is exploited numerically in Section 3, assuming $\beta=0.394$. This relationship could serve as a consistency check for future extensions to non-zero pressure gradient flows where $\kappa$ may vary.
3  Predictive Model and Insights
@@
It is remarked that this eddy corresponds to a closed loop in the sense that the leg of the eddy aligned with the wall is canceled by an equal and opposite image vortex pair, and is thus consistent with the vorticity kinematics. Using the Biot–Savart law to compute the induced velocity $u_{1}$ in Eq. (4), Figure 6 shows that such a simple attached eddy can reproduce the mean streamwise velocity accurately. Further, the second moments are also seen to be well-predicted as shown in Figure 6.
@@
Figure 6: Top figures show velocity moments. Symbols: Data; Lines: rectangular hairpin at $\theta=60^{\circ}$.. Bottom figures show eddy influence functions. Solid lines: Optimal eddy influence function. Dashed lines: rectangular hairpin at $\theta=60^{\circ}$. For the right figures, colors follow Figure 2.
It is noted that an extensive optimization was performed on different eddy shapes (e.g. triangular hairpin) and packet configurations. Yet, these did not yield improved predictions, particularly encountering difficulties in matching the log law slope while providing reasonable second moment and spectral behavior. The reasons behind this observation will be examined in detail in Section 5.
@@
The rectangular hairpin at an angle of $\theta=60^{\circ}$ is close to the optimal solution across different metrics and shows comparisons to the experimental measurements from Samie et al. (2018). As shown in Figures 7,8,9, the model remains remarkably accurate in predicting mean velocity, streamwise Reynolds stress and spectra over a range of Reynolds numbers.
We note that the influence functions were inferred from channel flow DNS at $Re_{\tau}\approx 5200$ (Lee & Moser, 2015), whereas the predictions in Figures 79 are compared against zero-pressure-gradient boundary layer measurements at $Re_{\tau}=6000$2000020000 (Samie et al., 2018). The fact that the same eddy template transfers successfully across flow geometries supports the universality of the log-layer structure, but this correspondence deserves further scrutiny, particularly for the Reynolds stress components and for the outer region where channel and boundary layer statistics diverge.
We also note that the model predictions align better with the data above the log layer than within it. This is not contradictory despite $p(h)\propto h^{-3}$ being questionable in the outer region: the rapid decay of $I_{1}(\eta)$ for $\eta>1$ ensures that the outer-region statistics are dominated by the largest eddies ($h\sim h_{\max}$), whose population amplitude is anchored by $\beta$ through the log-law calibration. The kernel structure naturally suppresses contributions from eddies smaller than the observation height, so the model works above the log layer primarily because of the influence-function shape rather than the population law.
@@
Figure 7: Mean velocity prediction using rectangular hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols:  Samie et al. (2018).
@@
Figure 8: Streamwise velocity fluctuation predictions using rectangular hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols:  Samie et al. (2018).
@@
Figure 9: Pre-multiplied streamwise energy spectra predictions at $y^{+}=1184$ using rectangular hairpin at $\theta=60^{\circ}$ for $Re_{\tau}=[6000,10000,14500,20000]$. Symbols:  Samie et al. (2018).
@@
The inverse problem strongly constrains what any viable attached eddy must do in a wall-parallel plane. In particular, the inferred mean influence function $I_{1}(y/h)$ exhibits (a) a near-constant region for $y/h\lesssim 1$ and (b) rapid decay for $y/h\gtrsim 1$, implying that eddies larger than the observation height dominate the log-region population integral. The rectangular hairpin + image realizes these features because the two legs and head generate partially canceling far-field contributions, while the inviscid image enforces wall compatibility and suppresses the leading-order tail that would otherwise make the planar influence integrals marginal (Appendix A). By contrast, a triangular hairpin produces a more monotone “ramp” in $I_{1}$ and does not reproduce the inferred plateau as cleanly (Figure 6). Experimentation with a pack of triangular hairpins with variable circulation strengths did not improve results beyond the rectangular hairpin which appears optimal.
@@
4  The Influence kernel
@@
$\mathcal{U}\!\left(\frac{x}{h},\frac{z}{h},\frac{y}{h}\right)=\mathcal{U}(\xi,\zeta,\eta),$
where $\mathcal{U}$ is the unit-eddy footprint expressed in dimensionless coordinates. Taking the Fourier transform and changing variables yields the exact scaling $\widehat{u}_{h}(k_{x},k_{z};y)\triangleq\,h^{2}\,\widehat{\mathcal{U}}(\kappa_{x},\kappa_{z};\eta),$ where $\widehat{\mathcal{U}}(\kappa_{x},\kappa_{z};\eta)\triangleq\int_{\mathbb{R}^{2}}\mathcal{U}(\xi,\zeta,\eta)\,e^{-i(\kappa_{x}\xi+\kappa_{z}\zeta)}\,d\xi\,d\zeta.$ Therefore
@@
where we define the Influence kernel as
@@
This identity is a concrete mathematical object that will help decode the energy spectrum in greater detail. In the following subsection, we will formulate an inverse problem to extract the shape of the ideal kernel.
It is noted that the spectral influence kernel is related to the one-point Reynolds stress kernel by $I_{11}(\eta)=\int_{0}^{\infty}I_{\phi}(\kappa_{x},\eta)\,d\kappa_{x},$ providing a consistency check between the spectral and one-point inverse problems.
4.1  Inverse Problem to Extract Ideal Influence Kernel
The inversion uses the attached-eddy spectral superposition in the form
$\phi_{xx}(k_{x};y)=\int_{h_{\min}}^{h_{\max}}\psi(k_{x}h,y/h)\,dh,$ (19)
where
$\psi(\kappa,\eta)=\beta CI_{\phi}(\kappa_{x},\eta),\qquad\kappa_{x}=k_{x}h,\qquad\eta=y/h.$
The numerical inversion solves for a gridded approximation to $\psi$. The kernel is represented on a tensor-product grid in $(\kappa,\eta)$, and the unknown vector is the row-major flattening of the $n_{\eta}\times n_{\kappa}$ kernel array:
$\bm{x}=\operatorname{vec}(\Psi)\in\mathbb{R}^{n_{\eta}n_{\kappa}},\qquad\Psi_{b,a}\approx\psi(\kappa_{a},\eta_{b}),$
with the index map $\ell(b,a)=(b-1)n_{\kappa}+a,\qquad 1\leq a\leq n_{\kappa},\quad 1\leq b\leq n_{\eta}.$ Hence $\bm{x}=[\Psi_{1,1},\Psi_{1,2},\ldots,\Psi_{1,n_{\kappa}},\Psi_{2,1},\ldots,\Psi_{n_{\eta},n_{\kappa}}]^{T}.$ Equation (19) is integrated over $h$ using $n_{h}$ geometrically spaced nodes $h_{1},\ldots,h_{n_{h}}\in[h_{\min},h_{\max}].$ The quadrature weights for a finite difference approximation are:
$w_{1}$ $=h_{2}-h_{1},$
$w_{q}$ $=\tfrac{1}{2}(h_{q+1}-h_{q-1}),\qquad 2\leq q\leq n_{h}-1,$
$w_{n_{h}}$ $=h_{n_{h}}-h_{n_{h}-1}.$
We then solve the non-negative regularized least-squares problem
$\min_{\bm{x}\geq 0}\frac{1}{2}\|\bm{A}\bm{x}-\bm{b}\|_{2}^{2}+\frac{\lambda_{\kappa}^{2}}{2}\|\bm{L}_{\kappa}\bm{x}\|_{2}^{2}+\frac{\lambda_{\eta}^{2}}{2}\|\bm{L}_{\eta}\bm{x}\|_{2}^{2}+\frac{\lambda_{0}^{2}}{2}\|\bm{L}_{0}\bm{x}\|_{2}^{2},$ (20)
Equivalently, this is written as an augmented non-negative least-squares problem
$\min_{\bm{x}\geq 0}\frac{1}{2}\|\widetilde{\bm{A}}\bm{x}-\widetilde{\bm{b}}\|_{2}^{2},$ (21)
with
$\widetilde{\bm{A}}=\begin{bmatrix}\bm{A}\\ \lambda_{\kappa}\bm{L}_{\kappa}\\ \lambda_{\eta}\bm{L}_{\eta}\\ \lambda_{0}\bm{L}_{0}\end{bmatrix},\qquad\widetilde{\bm{b}}=\begin{bmatrix}\bm{b}\\ \bm{0}\\ \bm{0}\\ \bm{0}\end{bmatrix}$.
For the reported results, $\widetilde{\bm{A}}\in\mathbb{R}^{28159\times 5712},\quad\operatorname{nnz}(\widetilde{\bm{A}})=1029731.$ and $\lambda_{\kappa}=10^{-1},\quad\lambda_{\eta}=4\times 10^{-1},\quad\lambda_{0}=10^{-5}.$ While these numbers appear arbitrary, the goal is to obtain insight into the shape of the kernel, and thus the primary metric is to obtain a smooth solution that does not overfit to measurement noise.
Appendix B provides more detail on the construction of these matrices. This is a more complicated inverse problem than the ones introduced earlier. Eq. (21) is solved by projected FISTA (Buccini et al., 2025). Figure 10 shows the reconstructed streamwise energy spectra (from the inverse solution) and how it compares to the data. The agreement is good, and regularization is shown to smooth the solution to a small degree, but not enough to impact any conclusions.
The spectral data used for this inversion are from the experimental measurements of Samie et al. (2018) at $Re_{\tau}=20000$. A brief sensitivity study on the regularization weights $(\lambda_{\kappa},\lambda_{\eta},\lambda_{0})$ confirmed that the qualitative shape of $I_{\phi}$, in particular the broad plateau at small $\eta$ and the localized hump at $\eta\sim O(1)$ is robust though finer features of the kernel near $\kappa_{x}\to 0$ are sensitive to the choice of $\lambda_{\kappa}$.
Figure 10: Data (symbols) and reconstructed (lines) streamwise energy spectra $\phi_{xx}^{+}(\kappa_{x})$ for a few wall normal locations
@@ 4.1  Inverse Problem to Extract Ideal Influence Kernel
Figure 11: Inferred streamwise energy Influence kernel $I_{\phi}(\kappa_{x},\eta)$ (normalized peak value). This figure highlights that the kernel has a broad, elevated region at small $\eta$ and small-to-moderate $\kappa$, and a more localized hump as $\eta$ increases. Also shown are contributions to $\phi_{xx}^{+}(k_{x}^{+})$ for $k_{x}^{+}=1\times 10^{-5}$ (cyan) and $k_{x}^{+}=1\times 10^{-4}$ (magenta) and $k_{x}^{+}=1\times 10^{-3}$ (red) at $y^{+}=1184$ for $Re_{\tau}=20000$.
@@
The extracted kernel $I_{\phi}(\kappa_{x},\eta)$, which is a function of the dimensionless wavenumber $\kappa_{x}=k_{x}h$ at fixed relative height $\eta=y/h$ and is shown in Figure 11. It is immediately clear that eddies of scale $h<y$ have negligible influence on the statistics.
@@
For any fixed $\eta$, typical scaling arguments yield a concentration of the influence $I_{\phi}(\kappa_{x},\eta)$ around $\kappa_{x}=O(1)$ because $\kappa_{x}$ is the only dimensionless streamwise wavenumber available to the unit eddy. Therefore, for a given size $h$ and fixed $y$, the dominant contribution of that eddy appears around
@@
equivalently $\lambda_{x}=2\pi/k_{x}=O(h)$. This explains the peak of the premultiplied spectra in Fig. 9.
The rectangular hairpin Attached Eddy model is able to recover a near-perfect linear behavior of the streamwise energy spectrum in the intermediate scale range as shown in Fig. 12. Below, we will give more concrete and quantitative explanations.
@@ 4.2  From single-eddy spectra to the attached-eddy spectrum
The $1/k_{x}$ scaling in the logarithmic region arises when $F(k_{x};y)$ varies slowly with $k_{x}$ over an intermediate band. This occurs when the $\kappa$-window $k_{x}h_{\min},k_{x}h_{\max}]$ covers the portion of the kernel $I_{\phi}(\kappa,\eta)$ that carries most of the contribution, and when the effective $\eta$ values encountered (through $\eta=y/h=yk_{x}/\kappa$) lie in the log-layer-relevant range. In that band,
@@
The above $1/k_{x}$” argument corresponds to the case that $F(k_{x};y)$ is approximately constant over an intermediate $k_{x}$ range. However the low-$k_{x}$ behavior is entirely controlled by how $F(k_{x};y)$ varies as $k_{x}\to 0$. There is no universality guarantee there: when $k_{x}$ is small, the $\kappa$-window and the $\eta$-trajectory can force the integral to sample the least robust part of the kernel.
@@
The overprediction of streamwise energy spectra by the Attached Eddy model in the high wavelength limit is clearly explained by Figures 14 and 15. For small $\kappa_{x}$, this trajectory spends substantial “arc length” in the region of small $\kappa$ and small $\eta$ and $I_{\phi}$ does not reduce to zero for small $\kappa_{x}$ as in Figure 11. Nevertheless, the overall shape of the kernel and the scale-by-scale contributions from the hairpin attached eddy appear impressive in relation to the inferred quantities.
Figure 14: Streamwise energy Influence kernel $I_{\phi}(\kappa_{x},\eta)$ (normalized peak value) for hairpin attached eddy at $Re_{\tau}=20000$.
Figure 15: Contribution to $F$ integral from different scales at $y^{+}=1184$ for $Re_{\tau}=20000$ for hairpin attached eddy.
@@
Physically, low $k_{x}$ corresponds to very long streamwise wavelengths. In the present attached-eddy framework with independent marks, the low-$k_{x}$ variance is built by incoherently summing contributions of the largest eddies. In a Poisson superposition, long-wavelength modes are fed by many large eddies with essentially random phases, so variances add without cancellation. Real wall turbulence exhibits organization, intermittency, and correlations at the largest scales; those effects can reduce energy in the very lowest modes relative to an independent-eddy superposition.  Baars & Marusic (2020a, b) offer more quantitative insights in the form of 3 different eddy types, and that attached eddies represent one of these types and contribute a large fraction of the energy only at an intermediate range of scales in the log layer. The latter point has been well argued by  Perry & Marušić (1995). The independent-mark hypothesis therefore tends to overpopulate the lowest wavenumbers unless correlations are explicitly modeled or packaged into larger marks. It is possible that the influence kernel $I_{\phi}$ can be used in an inverse problem setting to ‘design’ compositions of prototypical eddies to model desired spectral properties.
5  Why the rectangular hairpin is unusually predictive, and replacing it is difficult
The rectangular hairpin presented above is not just another convenient template that happens to fit well. In our optimization studies, we find that it occupies a singular corner of the design space. As will be discussed below, it concentrates the entire spanwise displacement at one height, which gives the exact step-shaped mean kernel required by the attached-eddy mean law, while its vertical legs remain spectrally broad and dominate the one-dimensional streamwise spectrum. In other words, the rectangular hairpin is unusually predictive because it realizes an unusually clean separation of duties. The top segment concentrates the entire spanwise displacement at one height and therefore gives the exact $I_{1}$ plateau. The legs carry essentially none of $I_{1}$, but because they are strongly localized they dominate most of $I_{\phi}$. The angle $\theta$ then tunes the spectral peak largely through phase, without spoiling the mean. Replacing the rectangular hairpin is hard because most alternatives spoil at least one of these properties, and many spoil two of them at once. This was a conclusion reached after a large number of optimization attempts. The mean behavior is definitively characterized by the discussion below.
Consider the symmetric three-segment hairpin family whose edge nodes are
$(0,0,-b_{0}/2)\rightarrow(\lambda,1,-b_{1}/2)\rightarrow(\lambda,1,b_{1}/2)\rightarrow(0,0,b_{0}/2)$
with the image included automatically in the discussion below. Figure 16 shows the geometry in the $(y,z)$ projection.
Figure 16: Straight-segment hairpin family and the corresponding exact $I_{1}(\eta)$ kernels.
Appendix C gives individual segment contributions.
Mean contribution
Summing the three segment contributions from Eq. (34) gives the exact single-hairpin mean kernel
$I_{1}^{\mathrm{hp}}(\eta)=\Gamma\,[b_{0}+(b_{1}-b_{0})\eta]H(1-\eta).$ (22)
This reduces to a plateau when $b_{1}=b_{0}$ and to a ramp when $b_{1}\neq b_{0}$. For the attached-eddy population, the mean profile becomes
$U^{+}(y^{+})=U_{\mathrm{ref}}^{+}-\beta C\Gamma F_{1}(y^{+}),$ (23)
with the exact population integral
$F_{1}(y^{+})=b_{0}\log\!\left(\dfrac{h_{\max}^{+}}{y^{+}}\right)+(b_{1}-b_{0})\left(1-\dfrac{y^{+}}{h_{\max}^{+}}\right),\ \ h_{\min}^{+}\leq y^{+}\leq h_{\max}^{+}.$ (24)
Thus, the rectangular hairpin $b_{1}=b_{0}=s$ collapses this to the exact logarithmic form
$F_{1}(y^{+})=s\log(h_{\max}^{+}/y^{+})\qquad(h_{\min}^{+}\leq y^{+}\leq h_{\max}^{+}).$
This is the main reason the square head remains so predictive: it gives the best possible morphology for the logarithmic mean.
Streamwise Kinetic Energy
For the rectangular hairpin with $0<\eta<1$, Appendix C provides exact expressions for the Fourier transform of the streamwise induced velocities. It is instructive to split the contributions from the head of the hairpin and the legs as $\hat{u}_{1}\triangleq\hat{u}_{h}+\hat{u}_{\ell}$. Then the spectral kernel can be reduced to an exact quadrature in $\kappa_{z}$:
$I_{\phi}^{hh}(\kappa_{x},\eta)$ $=\frac{1}{2\pi^{2}}\int_{0}^{\infty}\left|\widehat{u}_{h}(\kappa_{x},\kappa_{z};\eta)\right|^{2}\,d\kappa_{z},$ (25)
$I_{\phi}^{\ell\ell}(\kappa_{x},\eta)$ $=\frac{1}{2\pi^{2}}\int_{0}^{\infty}\left|\widehat{u}_{\ell}(\kappa_{x},\kappa_{z};\eta)\right|^{2}\,d\kappa_{z},$ (26)
$I_{\phi}^{h\ell}(\kappa_{x},\eta)$ $=\frac{1}{\pi^{2}}\operatorname{Re}\int_{0}^{\infty}\widehat{u}_{h}(\kappa_{x},\kappa_{z};\eta)\,\widehat{u}_{\ell}(\kappa_{x},\kappa_{z};\eta)^{*}\,d\kappa_{z}.$ (27)
For the representative rectangular hairpin at $\theta=60^{\circ}$, and $\eta=0.5$, these exact formulas give at $\kappa_{x}=0$
$I_{\phi}^{hh}(0,0.5)=2.74\times 10^{-2},\qquad I_{\phi}^{\ell\ell}(0,0.5)=8.36\times 10^{-2},\qquad I_{\phi}^{h\ell}(0,0.5)=4.82\times 10^{-2}.$
This reveals two things: a) The fact that the kernels (and thus $\phi_{xx}$) do not decay to zero at zero streamwise wavenumber; and b) even at $\kappa_{x}=0$, the head contributes only about $17\%$ of $I_{\phi}(0)$, the legs contribute about $53\%$, and the cross term contributes about $30\%$, even though the head contributes all of $I_{1}$. At the premultiplied peak ($\kappa_{x}\approx 1.55$) the split is approximately $8.5\%$, $71.4\%$, and $20.1\%$, and by $\kappa_{x}=5$ the legs carry about $98.1\%$ of the spectrum. This is the sharpest quantitative statement of the mean-variance duality: the head determines the mean kernel, but the legs determine most of the spectral energy.
Figure 17: Left: Exact square-hairpin decomposition of the single-eddy spectrum into head, leg, and cross contributions at $\eta=0.5$. Right: premultiplied spectrum for the same square geometry at several inclination angles. The mean kernel is fixed by the top segment, while the angle mainly tunes the spectral peak through the leg phase $e^{-i\kappa_{x}s\cot\theta}$.
As mentioned earlier, the difficulty of replacing the square head is precise rather than qualitative. Any replacement that spreads the top segment over a vertical band immediately destroys the exact plateau in $I_{1}$. Triangles, rounded tops, tilted tops, and arch-like heads all behave the same way.
The same replacement also changes the head contribution to $I_{\phi}$. Experiments show that downward spreading of the head not only spoils the mean plateau; it also increases the moderate-$\kappa_{x}$ head content because part of the head is brought closer to the observation plane. This is exactly the wrong direction if one wants to retain the clean head/leg role separation of the rectangular hairpin.
One might try to keep the top segment at a single height so that $I_{1}$ remains exact, and instead replace only the legs. This is a different failure mode. If the legs are broadened, bowed, packetized, or rounded in a way that gives them an effective spanwise width $w_{\ell}$, then the mean kernel is still unchanged to leading order because the leg contribution is small, but the spectrum is no longer the same. Even when the mean is preserved exactly, broadening the legs narrows the Cauchy spectrum and shifts the premultiplied peak to lower $\kappa_{x}$. In other words, keeping the right $I_{1}$ is not enough: the rectangular hairpin also succeeds because its legs are highly localized and therefore spectrally broad. That is the main reason the rectangular hairpin is unusually predictive: it is one of the very few simple filament templates in which the mean and spectral constraints are nearly decoupled.
5.1  Inferring and modeling the eddy density
It is useful to separate (i) the population law $(\beta,p(h),h_{\min},h_{\max})$ from (ii) the single-eddy template (geometry, inclination angle $\theta$, circulation scaling, and vortex-core regularization). Under these conventions, the only parameter used to anchor the data is the population amplitude $\beta$. It is indeed possible to improve the accuracy of the model by making the eddy density a function of the eddy size. A joint optimization run was performed for the eddy width, angle and $\beta(h^{+})$. The optimization for $Re_{\tau}=20000$ converged to $\theta=75^{\circ}$ and an aspect ratio (ratio of hairpin width and height) of 1.8, and an eddy density as shown in Figure 18, which imply that it is possible to get a near-perfect prediction. Keeping the eddy shape and angle fixed, and examining the behavior of the eddy density function, the following piecewise linear fit was found to be accurate
$C\beta(\bar{h})=\begin{cases}1.25+3.5\bar{h}&0\leq\bar{h}\leq 0.5,\\[11.99998pt] 3.0-2.5(\bar{h}-0.5)/0.5&0.5\leq\bar{h}\leq 1.\end{cases}$ (28)
where $\bar{h}=\frac{h^{+}-h_{\min}^{+}}{h_{\max}^{+}-h_{\min}^{+}}$.
Physically, the inferred $C\beta(\bar{h})$ peaks at mid-range eddy sizes ($\bar{h}\approx 0.5$) and decreases toward both the smallest and largest eddies. This is qualitatively consistent with the expectation that the largest eddies are geometrically constrained by the boundary layer thickness and may be less numerous than a pure $p(h)\propto h^{-3}$ law would predict, while the smallest eddies in the log layer interact with the buffer region where the attached-eddy hypothesis breaks down. We also note that the jointly optimized angle ($\theta=75^{\circ}$) and aspect ratio (1.8) differ from the values used in Section 3 ($\theta=60^{\circ}$, aspect ratio 1); this suggests that when $\beta(h)$ absorbs additional modeling flexibility, the optimal eddy shape shifts accordingly, and the two parameterizations should not be directly compared.
Figure 18: Left and middle: Joint optimal solution for eddy shape and density at $Re_{\tau}=20000$. Right: Inferred eddy density functions for $Re_{\tau}=[6000,10000,14500,20000]$ and piecewise linear approximation.
Using this analytical eddy density function, Figure 19 shows predictions at 4 different Reynolds numbers.
Figure 19: Improved predictions using analytical eddy density function $C\beta(h^{+})$ for $Re_{\tau}=[6000,10000,14500,20000]$.